HETAV PATEL

Grade 11,

How do you find the remainder of 27 raise to 27 divided by 1000 ?

How do you find the remainder of 27 raise to 27 divided by 1000 ?

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

We have
\frac{27^{27}}{1000}
\frac{(27^{9})^{3}}{(10)^{3}}
\frac{(30-3)^{9}}{10}
\frac{10k-3}{10}
k-\frac{3}{10}
k-1+1-\frac{3}{10}
(k-1)+(\frac{7}{10})
Remainder is 7.